OpenStudy (anonymous):

Complete the problem: 13 =3g - 2 (-5 + g)

OpenStudy (anonymous):

So what have you done? what's your approach?

OpenStudy (anonymous):

3g + 10 - 2g = 13 3g - 2g = 13 - 10 g = 3

OpenStudy (anonymous):

Please do not give out full answers

OpenStudy (anonymous):

All the teacher gave us was the problem and told us to complete it. I've been sitting here for like an hour trying to figure it out. I'm super bad at Algebra 2.